TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint8x2007x06x13x15x31

Powering of sector 78 - Report of 13.06.2007

Report of Wednesday 13.06.2007

  • ElQA
    • Tests in MBA during all the day.
  • RQ5.L8:
    • Calibration test performed by AB/PO. Test ongoing.
  • RD2.L8:
    • Ramped to 200A. Quench heater fired. Post-mortem data acquired properly. Test OK
    • Ramped to 1kA. Quench heater fired. Post-mortem data acquired properly. Test OK
  • RQF and RQD
    • Quench heater charge.
  • RD2.L8
    • Test 1 on the ML8 performed. D2 quenched when R4.L8B1 was at 2.2kA and R4.L8B2 was around 3kA ramping to 3.61kA. Test NOT OK.

Planning for Thursday 14.06.2007

AM:

  • ElQA on MB with instrumentation + QF & QD
  • RQ6.R7B1 and RQ6.R7B2: Repeat test
  • QPS IST RQF.A78 and RQD.A78: discharge of the heaters
  • RD2.L8 to nominal
  • RQ5.L8 and their correctors calibration by PO (UA)
PM:
  • 600 A circuits of L8: PCC (<20A) for all
  • 600 A circuits of L8: PIC2 for all
  • 600 A circuits of L8:PCS.1 for all
  • List of the 600A circuits released by AT/MEL and AT/ACR
    • RQTL12.L8B1
    • RQTF.A78B1
    • RQTF.A78B2
    • RSD1.A78B2
    • RQS.L8B1
    • RSD2.A78B1
    • RSF1.A78B1
    • RSF1.A78B2
    • RCO.A78B1
  • AB/PO Type tests on 60A converter

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2007-06-14 - BorisBellesia
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login